# Deep Learning and Machine Learning: Advancing Big Data Analytics and Management with Object-Oriented Programming

### v4

@@
Congratulations! After completing the introductory tutorials, you now have a solid understanding of the basics of deep learning. You’ve learned about fundamental neural network structures, training methods, and have gained some confidence in applying these techniques to simple tasks. However, the world of deep learning is vast [1], and what you’ve seen so far is just the beginning. There are more powerful models and algorithms that will help you tackle the complex problems we encounter in real-world applications.
In this section, we will guide you deeper into the exciting world of deep learning and help you further advance your skills. We will begin with a deep dive into advanced classifiers, discussing key architectures such as ResNet [2], which have become fundamental in fields like computer vision. You’ll learn how these models work, focusing on the key innovations, such as residual networks, that address issues like vanishing gradients and overfitting in deep networks.
Next, we will explore the world of object detection. Object detection goes beyond simply identifying whether an object exists in an image; it also involves locating where these objects are. This task is critical in fields like autonomous driving and smart surveillance. We’ll explore powerful models such as YOLO [3] and Faster R-CNN [4], equipping you with the tools needed to tackle complex visual tasks that require object localization [5].
@@
Finally, we’ll wrap up with a discussion on object-oriented programming (OOP) [6] in deep learning. As your projects grow in complexity, writing clean, modular, and reusable code becomes essential. We’ll explore how OOP principles can help you improve code readability, maintainability, and scalability, making your development process more efficient.
@@ 2.1.1 The Origins of Computing
The origins of computing can be traced back to the early 19th century, long before the existence of electronic computers. The British mathematician Charles Babbage designed the first mechanical computer, the “Difference Engine [7],” followed by the more ambitious “Analytical Engine [8].” Although these machines were never fully built during Babbage’s time, their conceptual frameworks laid the groundwork for modern computers.
The Analytical Engine was notable because it was programmable, thanks to input using punched cards [9]—a technique borrowed from the Jacquard loom. Ada Lovelace, Babbage’s collaborator, is widely recognized as the world’s first programmer. She wrote what is considered the first algorithm intended for a machine, understanding that computers could be used for more than just calculations. This visionary perspective paved the way for modern computing.
@@ 2.1.2 The Rise of Electronic Computers: World War II and Beyond
The development of computing accelerated during World War II. The demand for rapid and reliable calculations, especially for ballistics and code-breaking, led to the construction of early electronic computers. Machines like the British “Colossus [10]” and the American “ENIAC [11]” were built during this period. ENIAC was used to calculate artillery firing tables for the U.S. Army and solve complex mathematical problems.
One notable aspect of this era was the significant role women played in programming. Many of the early programmers, including those who worked on ENIAC, were women, such as Jean Jennings and Frances Bilas [12, 13]. They wrote programs using punched cards and physical switches, which were highly labor-intensive but a vital step toward modern software development.
@@ 2.1.3 Punched Cards and Early Programming
Punched cards [14] were an essential input method in early computing. Each punched card represented a line of code or a specific instruction. These cards were fed into the computer in the correct sequence, allowing the machine to execute complex tasks. While punched cards made it possible to store and process information, they also had limitations. Debugging errors involved physically examining and reordering cards, making the process slow and error-prone. Despite these challenges, this method of programming was foundational, and it shaped early computing practices.
@@ 2.2.1 The Birth of High-Level Languages
  • Fortran (1957) [15]: Designed by IBM for scientific and engineering applications, Fortran (Formula Translation) allowed complex mathematical formulas to be written and computed easily. It was a major step forward in making programming more accessible to scientists and engineers.
  • COBOL (1959) [16]: COBOL (Common Business-Oriented Language) was created for business and administrative tasks. Its syntax was designed to be close to English, enabling business professionals to understand code without specialized training.
@@ 2.2.2 Procedural Programming
By the 1960s and 1970s, the programming paradigm shifted toward procedural programming [17]. This approach involved writing software as a series of functions or procedures that performed specific tasks. Each function had a clear input-output mechanism, making the software more modular and easier to maintain. C, developed in 1972, became the most influential procedural programming language, as it allowed precise control over hardware while maintaining higher-level abstractions for logic and flow control.
@@ 2.2.3 The Development of Computer Hardware
During this time, the rapid advancement in computer hardware played a crucial role in shaping programming techniques. The invention of transistors and integrated circuits in the 1950s and 1960s allowed computers to become smaller, faster, and more reliable. By the 1970s, personal computers like the Apple II and the IBM PC emerged, bringing computing power to a much larger audience [18]. These advances enabled more sophisticated software development, as more memory and processing power became available for complex programs.
@@ 2.3.1 The Basic Concepts of Object-Oriented Programming
As computer systems grew more complex, a new programming paradigm called Object-Oriented Programming (OOP) emerged in the late 1970s and early 1980s [19, 17]. OOP’s core idea was to structure programs around “objects,” which encapsulate both data (attributes) and behavior (methods). Objects are instances of “classes,” which define a blueprint for creating similar objects [20].
@@ 2.3.2 Early Object-Oriented Languages
The first programming language to support OOP was Simula (1967) [21]. Simula introduced the concept of classes and objects to simulate real-world systems. This was followed by Smalltalk (1972) [22], which further refined the OOP model and is considered one of the first fully object-oriented programming languages. Smalltalk introduced the idea that ”everything is an object,” including primitive data types like numbers and strings.
@@ 2.3.3 The Popularity of C++ and Java
The popularity of OOP grew in the 1980s with the development of C++, which extended the C programming language with object-oriented features [23]. C++ became widely adopted due to its combination of procedural and object-oriented capabilities.
In the 1990s, Java was introduced, designed to be fully object-oriented and platform-independent. Java’s slogan, ”Write once, run anywhere,” reflected its ability to run on any device with a Java Virtual Machine (JVM) [24], further promoting the OOP model. Java has since become one of the most widely used programming languages for building enterprise applications.
@@ 2.4.2 Design Patterns
Design patterns [25] are proven solutions to common software design problems. They offer a standardized approach to solving problems that arise frequently during software development. The most commonly used design patterns include:
@@ 2.5.1 The Story of Python
Python was created in 1989 by Guido van Rossum with the intention of making a programming language that was easy to read and write [26]. Python was designed to bridge the gap between scripting and system programming languages by offering both power and simplicity. Python’s syntax emphasizes readability, which makes it an excellent choice for both beginners and experienced developers. Over the years, Python has grown into a versatile language used for web development, data science, machine learning, and more[27].
@@ 2.5.2 Object-Oriented Programming in Python
Python is a multi-paradigm language, meaning it supports multiple programming styles, including procedural, functional, and object-oriented programming. In Python, everything is an object, including primitive data types like integers and strings. Python allows developers to create their own classes and define attributes and methods for those classes[28].
@@ 2.6 Object-Oriented Programming in Artificial Intelligence
Object-Oriented Programming (OOP) is a powerful programming paradigm that structures code using objects and classes. This approach enables efficient and organized development, making complex systems like machine learning (ML)[5], deep learning (DL), large language models (LLM)[1], and data analytics more manageable, reusable, and scalable[29].
@@ 2.9.4 Attributes and Methods
[…]
@@ 2.10.1 Introduction
Encapsulation [30] is a fundamental concept in Object-Oriented Programming (OOP) that involves bundling the data (attributes) and methods (functions) that operate on the data within a single unit or class. It is one of the core principles of OOP, along with inheritance, polymorphism, and abstraction.
@@ 2.10.2 Access Modifiers (Private, Public, Protected)
Access modifiers [31] are special keywords or symbols used to control the visibility of class attributes and methods. In Python, access modifiers are achieved using naming conventions. Python does not enforce strict access control as in some other languages like Java or C++, but it does provide mechanisms to indicate how a member of a class should be accessed.
@@ 2.11.1 Introduction
Inheritance [32] is one of the fundamental concepts of Object-Oriented Programming (OOP) in Python. It allows one class to inherit attributes and methods from another class, promoting code reuse and establishing hierarchical relationships between classes. Inheritance helps to avoid redundancy and makes code more maintainable and scalable.
@@ 2.12.1 Introduction
Polymorphism [33] is one of the core concepts in Object-Oriented Programming (OOP). It refers to the ability of different classes to provide a unique implementation of the same method. The term polymorphism is derived from Greek words meaning ”many forms”. In Python, polymorphism allows methods to perform different functions based on the object that is calling them. This flexibility promotes code extensibility and reusability, as it allows objects of different classes to respond to the same message (method call) in different ways.
@@ 2.12.4 Dynamic Binding
Dynamic binding [34] is closely related to run-time polymorphism. It refers to the process by which the method that is to be executed is determined at runtime based on the actual object type. In Python, all method calls are dynamically bound because the method to be invoked is determined by the type of the object at runtime.
@@ 2.13.1 Introduction
Abstraction [35] is a fundamental concept in Object-Oriented Programming (OOP) that focuses on hiding the complex internal implementation details of a class or object and only exposing the necessary and relevant features to the user. This helps reduce complexity and allows the programmer to focus on interacting with the object rather than worrying about the inner workings of its methods or attributes.
@@ 2.14.2 Association
Association [36] is a general term used to define a relationship between two objects. In association, one object uses or interacts with another object without owning it. It’s a ”has-a” relationship, but the lifetime of the associated objects is independent. This is the most basic form of relationship between objects. For example, a teacher might be associated with multiple students, and a student might be associated with multiple teachers, but neither owns the other.
@@ 2.14.3 Aggregation
Aggregation [37] is a specialized form of association where one object contains another, but the contained object can exist independently of the container. This is still a ”has-a” relationship, but in this case, the lifetime of the contained object is not dependent on the container object. In other words, if the container object is destroyed, the contained object can still exist.
@@ 2.14.4 Composition
Composition [38] is a stronger form of aggregation where one object owns another object. In this case, the contained object cannot exist without the container. This implies a strict lifecycle dependency between the container and the contained object. When the container object is destroyed, the contained object is also destroyed.
@@ 2.15.1 Introduction
When working with object-oriented programming (OOP), particularly in Python, it is essential to follow design principles that help produce clean, efficient, and maintainable code. These principles guide developers in structuring code that is easy to understand, modify, and extend without introducing errors or unnecessary complexity. Some of the most widely accepted design principles are the SOLID principles, along with others like DRY and KISS [39]. Each of these principles ensures that your code remains organized and reduces the risk of introducing bugs during development or future changes.
@@ 2.15.2 SOLID Principles
The SOLID principles [40] are a set of five design principles that, when followed, enable developers to write better, more maintainable, and extensible object-oriented software. Let’s dive into each principle and explore it in the context of Python with clear examples.
@@
The Single Responsibility Principle (SRP) [41] states that a class should have only one reason to change, meaning that it should have only one responsibility. In simple terms, a class should do one thing and do it well. If a class has more than one responsibility, it becomes harder to maintain and modify without affecting other functionalities.
@@
The Open/Closed Principle (OCP) [30] dictates that software entities (such as classes, modules, or functions) should be open for extension but closed for modification. This means that you should be able to extend a class’s behavior without modifying its existing code, which reduces the risk of introducing bugs.
@@
The Liskov Substitution Principle (LSP) [42] states that objects of a subclass should be able to replace objects of the superclass without affecting the correctness of the program. This principle encourages us to ensure that subclasses behave in a way that their use in place of a parent class doesn’t break the functionality.
@@
The Interface Segregation Principle (ISP) [40] suggests that no client should be forced to depend on methods it does not use. In Python, where explicit interfaces are not used like in other languages, this principle can be applied by ensuring that classes don’t have unnecessary methods that are irrelevant to certain clients.
@@
The Dependency Inversion Principle (DIP) [43] states that high-level modules should not depend on low-level modules but both should depend on abstractions. This means that the details of implementation should depend on abstractions (interfaces) rather than the other way around.
@@ 2.15.3 DRY (Don’t Repeat Yourself)
The DRY principle [44] emphasizes avoiding repetition of code or logic. If you find yourself writing the same or similar code in multiple places, it’s better to abstract it into a function, class, or module that can be reused. This improves code maintainability and reduces the likelihood of errors.
@@ 2.15.4 KISS (Keep It Simple, Stupid)
The KISS principle [45] states that systems should be as simple as possible. Avoid overcomplicating solutions and aim to write clear, straightforward code that is easy to understand. Unnecessary complexity can lead to confusion, increased bugs, and difficulties when maintaining or extending the system.

### v5

@@
   Xinyuan Song
Emory Unversity
xsong30@emory.edu
@@
Congratulations! After completing the introductory tutorials, you now have a solid understanding of the basics of deep learning. You’ve learned about fundamental neural network structures, training methods, and have gained some confidence in applying these techniques to simple tasks. However, the world of deep learning is vast [niu2024large], and what you’ve seen so far is just the beginning. There are more powerful models and algorithms that will help you tackle the complex problems we encounter in real-world applications.
In this section, we will guide you deeper into the exciting world of deep learning and help you further advance your skills. We will begin with a deep dive into advanced classifiers, discussing key architectures such as ResNet [he2016identity], which have become fundamental in fields like computer vision. You’ll learn how these models work, focusing on the key innovations, such as residual networks, that address issues like vanishing gradients and overfitting in deep networks.
Next, we will explore the world of object detection. Object detection goes beyond simply identifying whether an object exists in an image; it also involves locating where these objects are. This task is critical in fields like autonomous driving and smart surveillance. We’ll explore powerful models such as YOLO [jiang2022review] and Faster R-CNN [jiang2017face], equipping you with the tools needed to tackle complex visual tasks that require object localization [Peng_2024].
@@
Finally, we’ll wrap up with a discussion on object-oriented programming (OOP) [rentsch1982object] in deep learning. As your projects grow in complexity, writing clean, modular, and reusable code becomes essential. We’ll explore how OOP principles can help you improve code readability, maintainability, and scalability, making your development process more efficient.
@@ 2.1.1 The Origins of Computing
The origins of computing can be traced back to the early 19th century, long before the existence of electronic computers. The British mathematician Charles Babbage designed the first mechanical computer, the “Difference Engine [ansell1997deleuze],” followed by the more ambitious “Analytical Engine [bromley1982charles].” Although these machines were never fully built during Babbage’s time, their conceptual frameworks laid the groundwork for modern computers.
The Analytical Engine was notable because it was programmable, thanks to input using punched cards [harlizius2017weaving]—a technique borrowed from the Jacquard loom. Ada Lovelace, Babbage’s collaborator, is widely recognized as the world’s first programmer. She wrote what is considered the first algorithm intended for a machine, understanding that computers could be used for more than just calculations. This visionary perspective paved the way for modern computing.
@@ 2.1.2 The Rise of Electronic Computers: World War II and Beyond
The development of computing accelerated during World War II. The demand for rapid and reliable calculations, especially for ballistics and code-breaking, led to the construction of early electronic computers. Machines like the British “Colossus [dell2014colossus]” and the American “ENIAC [weik1961eniac]” were built during this period. ENIAC was used to calculate artillery firing tables for the U.S. Army and solve complex mathematical problems.
One notable aspect of this era was the significant role women played in programming. Many of the early programmers, including those who worked on ENIAC, were women, such as Jean Jennings and Frances Bilas [todd2015jean, fritz1996women]. They wrote programs using punched cards and physical switches, which were highly labor-intensive but a vital step toward modern software development.
@@ 2.1.3 Punched Cards and Early Programming
Punched cards [casey1951punched] were an essential input method in early computing. Each punched card represented a line of code or a specific instruction. These cards were fed into the computer in the correct sequence, allowing the machine to execute complex tasks. While punched cards made it possible to store and process information, they also had limitations. Debugging errors involved physically examining and reordering cards, making the process slow and error-prone. Despite these challenges, this method of programming was foundational, and it shaped early computing practices.
@@ 2.2.1 The Birth of High-Level Languages
  • Fortran (1957) [backus1978history]: Designed by IBM for scientific and engineering applications, Fortran (Formula Translation) allowed complex mathematical formulas to be written and computed easily. It was a major step forward in making programming more accessible to scientists and engineers.
  • COBOL (1959) [sammet1978early]: COBOL (Common Business-Oriented Language) was created for business and administrative tasks. Its syntax was designed to be close to English, enabling business professionals to understand code without specialized training.
@@ 2.2.2 Procedural Programming
By the 1960s and 1970s, the programming paradigm shifted toward procedural programming [wegner1990concepts]. This approach involved writing software as a series of functions or procedures that performed specific tasks. Each function had a clear input-output mechanism, making the software more modular and easier to maintain. C, developed in 1972, became the most influential procedural programming language, as it allowed precise control over hardware while maintaining higher-level abstractions for logic and flow control.
@@ 2.2.3 The Development of Computer Hardware
During this time, the rapid advancement in computer hardware played a crucial role in shaping programming techniques. The invention of transistors and integrated circuits in the 1950s and 1960s allowed computers to become smaller, faster, and more reliable. By the 1970s, personal computers like the Apple II and the IBM PC emerged, bringing computing power to a much larger audience [hagedoorn2001strange]. These advances enabled more sophisticated software development, as more memory and processing power became available for complex programs.
@@ 2.3.1 The Basic Concepts of Object-Oriented Programming
As computer systems grew more complex, a new programming paradigm called Object-Oriented Programming (OOP) emerged in the late 1970s and early 1980s [nelson1990introduction, wegner1990concepts]. OOP’s core idea was to structure programs around “objects,” which encapsulate both data (attributes) and behavior (methods). Objects are instances of “classes,” which define a blueprint for creating similar objects [pokkunuri1989object].
@@ 2.3.2 Early Object-Oriented Languages
The first programming language to support OOP was Simula (1967) [nygaard1978development]. Simula introduced the concept of classes and objects to simulate real-world systems. This was followed by Smalltalk (1972) [kay1996early], which further refined the OOP model and is considered one of the first fully object-oriented programming languages. Smalltalk introduced the idea that ”everything is an object,” including primitive data types like numbers and strings.
@@ 2.3.3 The Popularity of C++ and Java
The popularity of OOP grew in the 1980s with the development of C++, which extended the C programming language with object-oriented features [stroustrup2007evolving]. C++ became widely adopted due to its combination of procedural and object-oriented capabilities.
In the 1990s, Java was introduced, designed to be fully object-oriented and platform-independent. Java’s slogan, ”Write once, run anywhere,” reflected its ability to run on any device with a Java Virtual Machine (JVM) [karakaya2008java], further promoting the OOP model. Java has since become one of the most widely used programming languages for building enterprise applications.
@@ 2.4.2 Design Patterns
Design patterns [coplien1998software] are proven solutions to common software design problems. They offer a standardized approach to solving problems that arise frequently during software development. The most commonly used design patterns include:
@@ 2.5.1 The Story of Python
Python was created in 1989 by Guido van Rossum with the intention of making a programming language that was easy to read and write [van2007python]. Python was designed to bridge the gap between scripting and system programming languages by offering both power and simplicity. Python’s syntax emphasizes readability, which makes it an excellent choice for both beginners and experienced developers. Over the years, Python has grown into a versatile language used for web development, data science, machine learning, and more[chen2024deeplearningmachinelearning].
@@ 2.5.2 Object-Oriented Programming in Python
Python is a multi-paradigm language, meaning it supports multiple programming styles, including procedural, functional, and object-oriented programming. In Python, everything is an object, including primitive data types like integers and strings. Python allows developers to create their own classes and define attributes and methods for those classes[peng2024securinglargelanguagemodels].
@@ 2.6 Object-Oriented Programming in Artificial Intelligence
Object-Oriented Programming (OOP) is a powerful programming paradigm that structures code using objects and classes. This approach enables efficient and organized development, making complex systems like machine learning (ML)[Peng_2024], deep learning (DL), large language models (LLM)[niu2024large], and data analytics more manageable, reusable, and scalable[peng2024deeplearningmachinelearning].
@@ 2.9.4 Attributes and Methods
Car (Class)car1 (Object) color = ”red” brand = ”Toyota” wheels = 4car2 (Object) color = ”blue” brand = ”Honda” wheels = 4
@@ 2.10.1 Introduction
Encapsulation [skoglund2003practical] is a fundamental concept in Object-Oriented Programming (OOP) that involves bundling the data (attributes) and methods (functions) that operate on the data within a single unit or class. It is one of the core principles of OOP, along with inheritance, polymorphism, and abstraction.
@@ 2.10.2 Access Modifiers (Private, Public, Protected)
Access modifiers [jackson2016objects] are special keywords or symbols used to control the visibility of class attributes and methods. In Python, access modifiers are achieved using naming conventions. Python does not enforce strict access control as in some other languages like Java or C++, but it does provide mechanisms to indicate how a member of a class should be accessed.
@@ 2.11.1 Introduction
Inheritance [cartwright1998empirical] is one of the fundamental concepts of Object-Oriented Programming (OOP) in Python. It allows one class to inherit attributes and methods from another class, promoting code reuse and establishing hierarchical relationships between classes. Inheritance helps to avoid redundancy and makes code more maintainable and scalable.
@@ 2.12.1 Introduction
Polymorphism [issariyakul2012review] is one of the core concepts in Object-Oriented Programming (OOP). It refers to the ability of different classes to provide a unique implementation of the same method. The term polymorphism is derived from Greek words meaning ”many forms”. In Python, polymorphism allows methods to perform different functions based on the object that is calling them. This flexibility promotes code extensibility and reusability, as it allows objects of different classes to respond to the same message (method call) in different ways.
@@ 2.12.4 Dynamic Binding
Dynamic binding [dewar2006safety] is closely related to run-time polymorphism. It refers to the process by which the method that is to be executed is determined at runtime based on the actual object type. In Python, all method calls are dynamically bound because the method to be invoked is determined by the type of the object at runtime.
@@ 2.13.1 Introduction
Abstraction [lieberman2006continuing] is a fundamental concept in Object-Oriented Programming (OOP) that focuses on hiding the complex internal implementation details of a class or object and only exposing the necessary and relevant features to the user. This helps reduce complexity and allows the programmer to focus on interacting with the object rather than worrying about the inner workings of its methods or attributes.
@@ 2.14.2 Association
Association [williams2002associative] is a general term used to define a relationship between two objects. In association, one object uses or interacts with another object without owning it. It’s a ”has-a” relationship, but the lifetime of the associated objects is independent. This is the most basic form of relationship between objects. For example, a teacher might be associated with multiple students, and a student might be associated with multiple teachers, but neither owns the other.
@@ 2.14.3 Aggregation
Aggregation [johnson1993refactoring] is a specialized form of association where one object contains another, but the contained object can exist independently of the container. This is still a ”has-a” relationship, but in this case, the lifetime of the contained object is not dependent on the container object. In other words, if the container object is destroyed, the contained object can still exist.
@@ 2.14.4 Composition
Composition [nierstrasz1995object] is a stronger form of aggregation where one object owns another object. In this case, the contained object cannot exist without the container. This implies a strict lifecycle dependency between the container and the contained object. When the container object is destroyed, the contained object is also destroyed.
@@ 2.15.1 Introduction
When working with object-oriented programming (OOP), particularly in Python, it is essential to follow design principles that help produce clean, efficient, and maintainable code. These principles guide developers in structuring code that is easy to understand, modify, and extend without introducing errors or unnecessary complexity. Some of the most widely accepted design principles are the SOLID principles, along with others like DRY and KISS [russell2024principles]. Each of these principles ensures that your code remains organized and reduces the risk of introducing bugs during development or future changes.
@@ 2.15.2 SOLID Principles
The SOLID principles [madasu2015solid] are a set of five design principles that, when followed, enable developers to write better, more maintainable, and extensible object-oriented software. Let’s dive into each principle and explore it in the context of Python with clear examples.
@@
The Single Responsibility Principle (SRP) [holderer2022single] states that a class should have only one reason to change, meaning that it should have only one responsibility. In simple terms, a class should do one thing and do it well. If a class has more than one responsibility, it becomes harder to maintain and modify without affecting other functionalities.
@@
The Open/Closed Principle (OCP) [skoglund2003practical] dictates that software entities (such as classes, modules, or functions) should be open for extension but closed for modification. This means that you should be able to extend a class’s behavior without modifying its existing code, which reduces the risk of introducing bugs.
@@
The Liskov Substitution Principle (LSP) [haoyu2012basic] states that objects of a subclass should be able to replace objects of the superclass without affecting the correctness of the program. This principle encourages us to ensure that subclasses behave in a way that their use in place of a parent class doesn’t break the functionality.
@@
The Interface Segregation Principle (ISP) [madasu2015solid] suggests that no client should be forced to depend on methods it does not use. In Python, where explicit interfaces are not used like in other languages, this principle can be applied by ensuring that classes don’t have unnecessary methods that are irrelevant to certain clients.
@@
The Dependency Inversion Principle (DIP) [thennakoon2022study] states that high-level modules should not depend on low-level modules but both should depend on abstractions. This means that the details of implementation should depend on abstractions (interfaces) rather than the other way around.
@@ 2.15.3 DRY (Don’t Repeat Yourself)
The DRY principle [keey2013drying] emphasizes avoiding repetition of code or logic. If you find yourself writing the same or similar code in multiple places, it’s better to abstract it into a function, class, or module that can be reused. This improves code maintainability and reduces the likelihood of errors.
@@ 2.15.4 KISS (Keep It Simple, Stupid)
The KISS principle [alwin2016kiss] states that systems should be as simple as possible. Avoid overcomplicating solutions and aim to write clear, straightforward code that is easy to understand. Unnecessary complexity can lead to confusion, increased bugs, and difficulties when maintaining or extending the system.